# Martin Boundary and the Nonlinear Multiplicative Heat Equation in Weak disorder

### v1

@@ Abstract
We study the nonlinear multiplicative heat equation on Dirichlet spaces with white in time noise under weak disorder. We show that positive invariant random fields with uniformly bounded second moments are in one-to-one correspondence with bounded positive harmonic functions. The proof combines a remote past pullback construction with a uniqueness argument that applies a contraction inspired by chaos expansion. As a consequence, the space of invariant measures inherits geometric structure from the Martin Boundary. We further establish a small-noise Gaussian fluctuation result within each harmonic sector and show that the long-time behavior of solutions is completely determined by the Martin boundary data of the initial condition. These results reveal a direct connection between stochastic PDE dynamics and boundary theory in potential analysis.
@@ 1 Introduction
Let $(M,d,\mu)$ be a metric measure space equipped with a Dirichlet form $\mathcal{E}$. Denote the heat semigroup and self-adjoint operator associated with $\mathcal{E}$ by $P_{t}$ and $\mathcal{L}$. We study the parabolic Anderson model (PAM), formally written as
@@
Here, $\beta>0$ is the inverse temperature parameter, $f:\mathbb{R}\to\mathbb{R}$ is a Lipschitz function satisfying $u~||f(u)-f(\tilde{u})|\leq L_{f}$|u-$\tilde${u}| and $f(0)=0$, and $\dot{W}=\dot{W}(t,x)$ is a centered Gaussian noise with covariance kernel $R$, (the exact definition will be given in section 2). We will assume $R$ satisfies the following assumption for this paper.
@@ 1.2 Acknowledgements
The author thanks Le Chen and Alexander Dunlap for helpful discussions which motivated this work, as well as Fabrice Baudoin for pointing out an error in a previous version.
@@ 2 Preliminaries
  • dzdz will denote integration with respect to the $\mu$ on $M$ in the variable $z$, and $f(z)dz$ for some measurable $f:M\to\mathbb{R}$ will denote the volume measure weighted by $f$.
We now introduce centered Gaussian noises with covariance kernel $R$ and rigorously define the notion of a solution. Any positive-definite function $R$ on $M$ defines an inner product defines an inner product by
@@
We denote the Hilbert space of space-time functions corresponding to $\braket{\cdot,\cdot}_{R}$ by $\mathcal{H}_{R}$. On a probability space $(\Omega,\mathcal{G},\mathbb{P})$, the noise $\dot{W}$ is defined as the isonormal Gaussian process[Nua95] over $\mathcal{H}_{R}$ such that $\mathcal{G}=\sigma(\dot{W})$. We will not denote its $R$-dependence as there is no concern for confusion.
@@ 3.1 Existence by Backward Time Shifts
where $S_{\tau}$ denotes the semigroup corresponding to (1) acting on probability measures over $C_{\mathrm{loc}}(\mathbb{H}^{d})$.
@@ Appendix A Appendix: Connections to Martin Boundary
    (Equivariance of the dynamics.) For every $g\in\mathrm{Isom}(\mathbb{H}^{d})$ and $t\geq 0$,
@@
    (Equivariance of invariant measures.) If $\nu$ is an invariant measure for the Markov semigroup $\{\mathcal{S}_{t}^{W}\}_{t\geq 0}$ (i.e. $u_{0}\sim\nu\implies\mathcal{S}_{t}^{W}(u_{0})\sim\nu$ for all $t\geq 0$), then $g_{\#}\nu$ is also invariant for every $g\in\mathrm{Isom}(\mathbb{H}^{d})$.
@@
    (Equivariance of invariant fields.) If $Z$ is an invariant random field (i.e. $Z\stackrel{{\scriptstyle\mathrm{law}}}{{=}}\mathcal{S}_{t}^{W}(Z)$ for all $t\geq 0$ with $Z$ independent of $W$), then $g\cdot Z$ is also invariant for every $g\in\mathrm{Isom}(\mathbb{H}^{d})$.

### v2

# Martin Boundary and the Nonlinear Multiplicative Stochastic Heat Equation in Weak Disorder

@@ Abstract
We study the nonlinear multiplicative stochastic heat equation on Dirichlet spaces with white in time noise under weak disorder. We show that positive invariant random fields with uniformly bounded second moments are in one-to-one correspondence with bounded positive harmonic functions. The proof combines a remote past pullback construction with a uniqueness argument that applies a contraction inspired by chaos expansion. As a consequence, the space of invariant measures inherits geometric structure from the Martin Boundary. We further establish a small-noise Gaussian fluctuation result within each harmonic sector and show that the long-time behavior of solutions is completely determined by the Martin boundary data of the initial condition. These results reveal a direct connection between stochastic PDE dynamics and boundary theory in potential analysis.
@@ 1 Introduction
Let $(M,d,\mu)$ be a metric measure space equipped with a symmetric Dirichlet form $\mathcal{E}$. Denote the heat semigroup and $L^{2}(M)$-self-adjoint operator associated with $\mathcal{E}$ by $P_{t}$ and $\mathcal{L}$, respectively. We study the multiplicative stochastic heat equation, formally written as
@@
Here, $\beta>0$ is the inverse temperature parameter, $f:\mathbb{R}\to\mathbb{R}$ is a Lipschitz function satisfying $u~||f(u)-f(\tilde{u})|\leq L_{f}$|u-$\tilde${u}| and $f(0)=0$, and $\dot{W}=\dot{W}(t,x)$ is a centered Gaussian noise with covariance kernel $R$ (the exact definition will be given in section 2). We will assume $R$ satisfies the following assumption for this paper.
@@ 1.2 Acknowledgements
The author thanks Le Chen and Alexander Dunlap for helpful discussions which motivated this work, as well as Fabrice Baudoin for pointing out an error in a previous version. This work was supported by a research grant (VIL73729) from Villum Fonden.
@@ 2 Preliminaries
  • dzdz will denote integration with respect to the $\mu$ on $M$ in the variable $z$.
We now introduce centered Gaussian noises with covariance kernel $R$ and rigorously define the notion of a solution. Any positive-definite function $R$ on $M$ defines an inner product by
@@
We denote the Hilbert space of space-time functions corresponding to $\braket{\cdot,\cdot}_{R}$ by $\mathcal{H}_{R}$. On a probability space $(\Omega,\mathcal{G},\P)$, the noise $\dot{W}$ is defined as the isonormal Gaussian process[Nua95] over $\mathcal{H}_{R}$ such that $\mathcal{G}=\sigma(\dot{W})$. We will not denote its $R$-dependence as there is no concern for confusion.
@@ 3.1 Existence by Backward Time Shifts
where $S_{\tau}$ denotes the semigroup corresponding to (1) acting on probability measures over $C_{\mathrm{loc}}(M)$.
@@ Appendix A Appendix: Connections to Martin Boundary
    (Equivariance of the dynamics.) For every $g\in\mathcal{G}$ and $t\geq 0$,
@@
    (Equivariance of invariant measures.) If $\nu$ is an invariant measure for the Markov semigroup $\{\mathcal{S}_{t}^{W}\}_{t\geq 0}$ (i.e. $u_{0}\sim\nu\implies\mathcal{S}_{t}^{W}(u_{0})\sim\nu$ for all $t\geq 0$), then $g_{\#}\nu$ is also invariant for every $g\in\mathcal{G}$.
@@
    (Equivariance of invariant fields.) If $Z$ is an invariant random field (i.e. $Z\stackrel{{\scriptstyle\mathrm{law}}}{{=}}\mathcal{S}_{t}^{W}(Z)$ for all $t\geq 0$ with $Z$ independent of $W$), then $g\cdot Z$ is also invariant for every $g\in\mathcal{G}$.